Question 4: A firm has the following demand: P = 80-0.04Q a. Write down the total revenue function b. At what output level, total revenue will be maximized c. If Q = 200, what is the total revenue? d. Is the demand elastic or inelastic at Q = 200
Answer #1

a) TR = PQ = 80Q - 0.04Q2

b) TR is maximized where d(TR)/dQ = 0

80 - 0.08Q = 0

Q = 1000

c) when Q = 200

TR = 14400

d) Price elasticity = dQ/dP x P/Q

P = 80 - 0.04 x 200 = 72

dQ/dP = -25

PE = -25 x 72/200 = - 9 (demand is elastic as absolute value > 1)